# Two-sample Comparison through Additive Tree Models for Density Ratios

### 11 Mar 2026 v3

@@ 4.1 Two-dimensional Simulations
We simulate data sets using the following three scenarios which represent different types of distributional differences between $X_{0}$ and $X_{1}$:
1. Global shift
2. Local shift
where $\delta=(0,1)^{\prime}$, and $\mu_{k}$ and $\Sigma_{k}$ ($k=1,\dots,5$) are provided in Supplementary Materials C.
3. Local dispersion difference
where $\Delta=\mathrm{diag}(0.36,1)$, and $\mu_{k}$ and $\Sigma_{k}$ ($k=1,\dots,3$) are provided in Supplementary Materials C.
The simulated data sets are visualized in Figure 2. The first scenario represents a global-scale difference between the two samples, whereas the second and third scenarios involve more localized discrepancies.
@@ 3. Local dispersion difference
Figure 3 compares the estimated log-density ratios in the local shift scenario with the balanced sample sizes ($n_{0}=n_{1}=5000$). The estimated functions using the proposed methods successfully capture the local difference structure, and importantly, the log-density ratios are significantly positive or negative for many observations included in the mixture component that has the local shift difference. For example, in the plot of the 2.5% quantiles, we find observations with yellow/red colors, indicating that the log-density ratios are significantly larger than 0, mostly in the corresponding mixture component. Similar patterns are obtained for the other scenarios and sample size settings, and we provide their visualizations in Supplementary Materials D.
Figure 3: The estimated log-density ratios obtained from the algorithms considered in the two-dimensional examples for the local shift scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.
@@ 4.2 Higher-dimensional Settings
1. Location Shift
$Z_{0}\sim\frac{4}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),\ Z_{1}\sim\frac{4}{5}N(\mu_{1}+\delta,\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),$
where $\delta=(1,0,0,0)^{T}$. The parameters $\mu_{k}$ and $\Sigma_{k}$ ($k=1,2$) are provided in Supplementary Materials C.
2. Dispersion Difference
$Z_{0}\sim\frac{4}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),\ Z_{1}\sim\frac{4}{5}N(\mu_{1},\Delta\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),$
where $\Delta=\mathrm{diag}(0.5,1,1,1)$. The parameters $\mu_{k}$ and $\Sigma_{k}$ ($k=1,2$) are provided in Supplementary Materials C.
The orthonormal matrix $\Lambda$ is generated using the randortho function in the R package pracma (Borchers, 2023).
@@ 2. Dispersion Difference
5 Assessing the Quality of Synthesized Microbiome Compositions
@@ 6 Conclusion
We have introduced a new approach to learning the density ratio between two i.i.d. samples using the class of additive tree models and a new loss function called the balancing loss, which was shown to be highly effective in learning density ratios. Not only does it provide effective boosting algorithms to attain functional (point) estimates, but it also provides a convenient means for generalized Bayesian inference on the density ratio, attaining uncertainty quantification, which can be critical in challenging problems with limited sample size or high-dimensionality. The training algorithms we presented, both for boosting and for generalized Bayesian learning, can be introduced with modifications of their counterparts widely used in supervised learning based on this model class, and so their implementation would involve only minor adjustments to existing algorithms for boosting and BART models.
@@ Acknowledgment
This research is funded in part by NIGMS grant R01-GM135440 and NSF grant DMS-2152999. Part of the research was completed when LM and YX were at Duke University. NA’s research is supported by the Japan Society for the Promotion of Science KAKENHI (grant numbers: 25K16618, 24H00142, and 25H00546). Portions of the manuscript were prepared with the assistance of ChatGPT (from GPT-4.1 to 5.4) for English-language editing.
@@ A.2 Derivation of the regularized prior for the node parameters
the distribution of $\log w$ is approximated by $N(0,\lambda_{0}^{-1})$.
@@ B.2 Updating trees
$\alpha=\underbrace{\frac{p(T^{*}_{k})}{p(T_{k})}}_{\text{Prior}}\underbrace{\frac{p(T^{*}_{k}\to T_{k})}{p(T_{k}\to T^{*}_{k})}}_{\text{Transition}}\underbrace{\frac{p(\mathrm{data}\mid T_{k},\{w_{-k}(x_{i})\}^{n}_{i=1})}{p(\mathrm{data}\mid T^{*}_{k},\{w_{-k}(x_{i})\}^{n}_{i=1})}}_{\text{Loss-based likelihood}}$
@@ 3. Local dispersion difference
$\mu_{1}=(1.9,-7.2)^{\prime}$, $\mu_{2}=(-5.7,5.3)^{\prime}$, $\mu_{3}=(-2.3,-1.5)^{\prime}$, $\mu_{4}=(7.5,-3.1)^{\prime}$, $\mu_{5}=(3.1,9.5)^{\prime}$, $(\Sigma_{1}(1,1),\Sigma_{1}(1,2),\Sigma_{1}(2,2))=(1.0,-0.4,0.8)$,
$(\Sigma_{2}(1,1),\Sigma_{2}(1,2),\Sigma_{2}(2,2))=(1.3,0.7,2.7)$, $(\Sigma_{3}(1,1),\Sigma_{3}(1,2),\Sigma_{3}(2,2))=(1.0,0,3.0)$,
$(\Sigma_{4}(1,1),\Sigma_{4}(1,2),\Sigma_{4}(2,2))=(2.9,0,1.1)$, and $(\Sigma_{5}(1,1),\Sigma_{5}(1,2),\Sigma_{5}(2,2))=(2.4,-0.9,1.6)$.
C.2 20-dimensional dimensional experiments
@@ C.3 Notes on the implementation of the density ratio estimators
In the proposed boosting algorithms, we prune leaf nodes that contain observations from only one group to avoid division by zero in computing the optimal values of leaf node parameters.
@@ D.3 Alternative simulation scenarios with 20-dimensional structures
As a result, the marginal distributions are right-skewed and have a large probability around 0, and the distributional structure is similar to the high-dimensional abundance data, as illustrated in Figure 12. The estimation results based on 50 simulated data sets are summarized in Table 3, and it shows that our proposed methods (BS, FS, and BAT) still estimate the two-sample difference more accurately.
@@ Appendix E Additional real data analysis results
Train v.s. Generated: Posterior mean of DRE
Train v.s. Generated: Lower 2.5% MCMC quantile of DRE
Train v.s. Generated: Upper 97.5% MCMC quantile of DRE
@@
Test v.s. Generated: Posterior mean of DRE
Test v.s. Generated: Lower 2.5% MCMC quantile of DRE
Test v.s. Generated: Upper 97.5% MCMC quantile of DRE

### 22 Apr 2026 v4

@@ 4.1 Two-dimensional Simulations
We simulate data sets using the following three scenarios which represent different types of distributional differences between $X_{0}$ and $X_{1}$ visualized in Figure 2. The first scenario represents a global-scale difference between the two samples, whereas the second and third scenarios involve more localized discrepancies. The details are provided in Supplementary Materials C
@@
Figure 3 compares the estimated log-density ratios in the local shift scenario with the balanced sample sizes ($n_{0}=n_{1}=5000$). The estimated functions using the proposed methods successfully capture the local difference structure, and importantly, the log-density ratios are significantly positive or negative for many observations included in the mixture component with the local shift difference. For example, in the plot of the 2.5% quantiles, we find observations with yellow/red colors, indicating that the log-density ratios are significantly larger than 0, mostly in the corresponding mixture component. Similar patterns are obtained for the other scenarios and sample size settings, as detailed in Supplementary Materials D.
Figure 3: The estimated log-density ratios obtained from the algorithms considered in the two-dimensional examples for the local shift scenario ($n_{0}=n_{1}=5000$). For Bayesian additive trees, the pointwise 2.5% quantiles and 97.5% quantiles of the log-density ratios are also displayed. The difference between FS and GB is minimal, so we show only the results for GB.
@@ 4.2 Higher-dimensional Settings
  1. 1.
    “Location Shift”: $Z_{0}\sim\frac{4}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),\ Z_{1}\sim\frac{4}{5}N(\mu_{1}+\delta,\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2})$, where $\delta=(1,0,0,0)^{T}$.
  2. 2.
    “Dispersion Difference”: $Z_{0}\sim\frac{4}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),\ Z_{1}\sim\frac{4}{5}N(\mu_{1},\Delta\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2})$, where $\Delta=\mathrm{diag}(0.5,1,1,1)$.
The parameters ($\mu_{k}$ and $\Sigma_{k}$ for $k=1,2$) are provided in Supplementary Materials C. The orthonormal matrix $\Lambda$ is generated using the randortho function in the R package pracma (Borchers, 2023).
@@
5 Assessing Synthesized Microbiome Compositions
@@ 6 Conclusion
We have introduced a new approach to learning the density ratio between two i.i.d. samples using the class of additive tree models and a new loss function called the balancing loss. Not only does it provide effective boosting algorithms to attain functional (point) estimates, but it also provides a convenient means for attaining uncertainty quantification, which can be critical in challenging problems with limited sample size or high-dimensionality.
@@ Acknowledgment
This research is funded in part by NIGMS grant R01-GM135440 and NSF grant DMS-2152999. Part of the research was completed when LM and YX were at Duke University. NA’s research is supported by the Japan Society for the Promotion of Science KAKENHI (grant numbers: 25K16618, 24H00142, and 25H00546). Portions of the manuscript were prepared with the assistance of ChatGPT (from GPT-4.1 to 5.4) for English-language editing and coding.
@@ A.2 Derivation of the regularized prior for the node parameters
the distribution of $\log w(x)$ for fixed $x$ is approximated by $N(0,\lambda_{0}^{-1})$.
@@ B.2 Updating trees
$\alpha=\underbrace{\frac{p(T^{*}_{k})}{p(T_{k})}}_{\text{Prior}}\underbrace{\frac{p(T^{*}_{k}\to T_{k})}{p(T_{k}\to T^{*}_{k})}}_{\text{Transition}}\underbrace{\frac{p(\mathrm{data}\mid T^{*}_{k},\{w_{-k}(x_{i})\}^{n}_{i=1})}{p(\mathrm{data}\mid T_{k},\{w_{-k}(x_{i})\}^{n}_{i=1})}}_{\text{Loss-based likelihood}}$
@@ Appendix C Details of experiments
1. Global shift
2. Local shift
where $\delta=(0,1)^{\prime}$.
3. Local dispersion difference
where $\Delta=\mathrm{diag}(0.36,1)$.
@@ 3. Local dispersion difference
$\mu_{1}=(1.9,-7.2)^{\prime}$, $\mu_{2}=(-2.3,-1.5)^{\prime}$, $\mu_{3}=(7.5,-3.1)^{\prime}$, $(\Sigma_{1}(1,1),\Sigma_{1}(1,2),\Sigma_{1}(2,2))=(1.0,-0.4,0.8)$, $(\Sigma_{2}(1,1),\Sigma_{2}(1,2),\Sigma_{2}(2,2))=(1.0,0,3.0)$, and $(\Sigma_{3}(1,1),\Sigma_{3}(1,2),\Sigma_{3}(2,2))=(2.9,0,1.1)$.
C.2 20-dimensional experiments
@@ C.3 Notes on the implementation of the density ratio estimators
For boosting, we assess both the FS and the GB algorithms. The learning rate $\nu$ is set to 0.01, and the maximum resolution of each tree is set to 4. The optimal number of trees is selected with the 5-fold cross-validation with the maximum number allowed set to 1000. In both types of the boosting algorithms, we introduce 31 equally spaced cut points in each dimension. For a fair comparison, AdaBoost is implemented under the same tuning parameters. For the generalized Bayesian additive tree model, we set the size of the ensemble to 200, following the standard implementation of the BART (Chipman et al., 2010). The MCMC sampler is initialized with trees containing only root nodes, and the size of the burn-in and the posterior sample is set to 2000 and 1000, respectively.
Additionally, in the boosting algorithms, we prune leaf nodes that contain observations from only one group to avoid division by zero in computing the optimal values of leaf node parameters.
@@
We found that the compared methods can output infinite values. Such values are omitted in computing the means and the standard errors of the results. We note that such phenomena were not observed for the proposed methods.
@@ D.3 Alternative simulation scenarios with 20-dimensional structures
As a result, the marginal distributions are right-skewed and have a large probability around 0, and the distributional structure is similar to the high-dimensional abundance data, as illustrated in Figure 12. The estimation results based on 50 simulated data sets are summarized in Table 3, and it shows that our proposed methods (GB, FS, and BAT) still estimate the two-sample difference more accurately.
@@ Appendix E Additional real data analysis results
Train vs. Generated: Posterior mean of DRE
Train vs. Generated: Lower 2.5% MCMC quantile of DRE
Train vs. Generated: Upper 97.5% MCMC quantile of DRE
@@
Test vs. Generated: Posterior mean of DRE
Test vs. Generated: Lower 2.5% MCMC quantile of DRE
Test vs. Generated: Upper 97.5% MCMC quantile of DRE